OpenStudy (anonymous):

simplify the expression? a) (g^2h)^3 x (g^-2h^4)^-6

OpenStudy (anonymous):

Hint... You're supposed to know these two basic rules... \[\Large (g^2h)^3 \cdot (g^{-2}h^4)^{-6}\] \[\LARGE\boxed{\left(a^n \right)^m=a^{n\cdot m} \quad \quad \& \quad \quad a^m\cdot a^n=a^{m+n}}\]

OpenStudy (callisto):

\[ (g^2h)^3 x (g^{-2}h^4)^{-6}\] \[= (g^2h)^3 / (g^{-2}h^4)^6\] \[= g^{(2\times3)} h^3 / [ g^{(-2\times6)} h^{4\times6}) ]\] \[= g^6h^3 / g^{-12}h^{24}\] \[= g^6 g^{12}h^3/h^{24}\] Can you simplify it from here?